A pharmacist has 30% solution of a certain drug and a 12% solution of the same drug. How much of each solution should she mix, in order to make 6 liters of a 24% solution of the drug?
Answer by checkley77(12844)

.30X+.12(6-X)=.24*6
.30X+.72-.12X=1.44
.18X=1.44-.72
.18X=.72
X=.72/.18
X=4 LITERS OF THE 30% SOLUTION IS USED.
PROOF:
.30*4+.12(6-4)=1.44
1.2+.12*2=1.44
1.2+.24=1.44
1.44=1.44